Using traditional methods, it takes 99 hours to receive a basic driving license. A new license training method using Computer Aided Instruction (CAI) has been proposed. A researcher used the technique with 170 students and observed that they had a mean of 100 hours. Assume the variance is known to be 36. A level of significance of 0.05 will be used to determine if the technique performs differently than the traditional method. Find the value of the test statistic. Round your answer to 2 decimal places.

Enter the value of the test statistic.
Answer #1

H 997 +m 99 n=120, र =००, 2 - 36, =0-05 a) ema 2 R -99 | ป า 9 217 न In